OpenStudy (anonymous):

Solve the equation. 8x^3-3=0 Would it be like this? 8x^3=3 and then you have to cube both sides?

OpenStudy (anonymous):

First add 3 to both sides 8x^3 = 3 Now you can either divide or cube root, lets divide (8x^3) / 8 = 3/8 x^3 = 3/8 now take the cube root of 3/8

OpenStudy (anonymous):

Thank you!